OpenStudy (chelseatheruler):

how would you rationalize 3 over sqrt(5)

OpenStudy (ranga):

multiply top and bottom by sqrt(5).

OpenStudy (chelseatheruler):

What do you mean? @ranga I don't really understand this.

OpenStudy (ranga):

\[\frac{ 3 }{ \sqrt{5} } = \frac{ 3 \times \sqrt{5} }{ \sqrt{5} \times \sqrt{5} } = ?\]